p9x + 23/2 Evaluate Evaluate ſistemi det dr V
Answer #1

S(x+20h) dx = 5 (5x + x) dx S (ak tx) dx (232 => 3 [934– 1] + [(9124] = {[(332)%– 1] +4x80 Tegheria (1 263 + 40 = 172